A Math class has only 4 students (A, B, C and D) and they independently attend the class. The probability of them attending is 0.5, 0.6, 0.7, and 0.75, respectively. Find the 1 probability that the total number of attendance on 2 consecutive days is exactly 4?
Answer #1

Ans:

Probabiliity of attending for exactly 4 students on any day=0.5*0.6*0.7*0.75=0.1575

Now,

Probability of attending for exactly students on 2 consecutive days

=P(exactly 4 on day 1)*P(exactly 4 on day 2)

=0.1575*0.1575

=0.0248